Circular No. 8127

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 H4 M. Bezpalko, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the discovery by LINEAR of a comet with a tail in p.a. 270^{o} on images taken on Apr. 29.3 UT (discovery observation below). Following posting on the NEO Confirmation Page, other CCD observers have also reported the object as cometary, including G. J. Garradd (Tamworth, N.S.W., 0.45-m reflector; slightly diffuse on most images taken on Apr. 30.6), J. E. McGaha (Tucson, AZ, 0.30-m reflector; faint coma of size 5′′ × 10′′ and m_{1} = 17.7–17.9, aligned north-south, with uniform brightness and no apparent nuclear condensation or core on May 2.2), and J. G. Ries (McDonald Observatory, 0.76-m reflector; 20′′ tail pointing slightly south of west on May 2.3; m_{1} = 17.7–18.0).

2003 UT  α_{2000}  δ_{2000}  m_{1}
Apr. 29.32742  16^{h}06^{m}33.^{s}56  −19^{o}09′49′′.0  19.0

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2003-J14. These elements indicate that the comet passed 0.4 AU from Jupiter in Aug. 2000.

T = 2003 May 13.417 TT  ω = 10.^{o}054
e = 0.47584  Ω = 226.635 }2000.0
q = 1.68842 AU  i = 17.746
a = 3.22117 AU n^{o} = 0.170484 P = 5.78 years

V4745 SAGITTARII Further photoelectric photometry by A. C. Gilmore at Mt. John (cf.

IAUC 8123; uncertainties 0.02 mag or less in all colors): Apr. 28.635 UT, V = 9.31, U − B = −0.84, B − V = −0.09, V − R = +0.55, V − I = +0.32, airmass 1.09; 29.506, 9.37, −0.88, −0.08, +0.54, +0.28, 1.33.

CCD photometry by J. D. West, Mulvane, KS: Apr. 27.46 UT, V = 9.65, R_{c} = 8.90; 29.439, V = 9.22; 30.425, V = 9.55.

Concerning his magnitudes on IAUC 8123, M. Yamamoto adds that he used a green (PO0) filter whose reponse approaches that of the V band.

CORRIGENDA On IAUC 8089, second item, lines 10 and 11, the designations should

read SN 2003bo and SN 2003bp.

2003 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green